OpenStudy (anonymous):

if the equation 2x^3 +kx - 4 =0 and 6x^4 + 3kx^2 + 2 =0 have a common root k equals ?

OpenStudy (anonymous):

@Callisto @experimentX @badi

OpenStudy (callisto):

Can I do it in a very.... stupid way? 2x^3 +kx - 4 =0 kx = 4-2x^3 6x^4 + 3kx^2 + 2 =0 6x^4 + 3(kx)x + 2 =0 6x^4 + 3x(4-2x^3) +2 =0 6x^4 + 12x - 6x^4 +2 =0 Solve x

OpenStudy (callisto):

Then, sub. x into the equation and find k

OpenStudy (anonymous):

@Callisto by ur method i got k=71 but the answer should bek=-433/18

OpenStudy (callisto):

Nope... I got your answer :|

OpenStudy (anonymous):

I also got -433/18 for k. x=-1/6 and if you put this into k=(4-2x^3)/x you get the k you want.

OpenStudy (callisto):

2x^3 +kx - 4 =0 kx = 4-2x^3 6x^4 + 3kx^2 + 2 =0 6x^4 + 3(kx)x + 2 =0 6x^4 + 3x(4-2x^3) +2 =0 6x^4 + 12x - 6x^4 +2 =0 12x = -2 x= -1/6 kx = 4-2x^3 k = (4-2x^3) / x = (4-2(-1/6)^3) /(-1/6) = -433/18

OpenStudy (shubhamsrg):

let common root be c so we have 2c^3 + kc -4 =0 {multiply this by 3c} =>6c^4 + 3kc^2 - 12c=0 and 2nd is 6c^4 + 3kc^2 + 2=0 =>c=-1/6 thus,on substituting,, you should get your ans,,

OpenStudy (anonymous):

@Callisto sorry that my calculation was wrong

OpenStudy (callisto):

Never mind :)

OpenStudy (anonymous):

anywhere thanzz a lot....

OpenStudy (anonymous):

x=-1/6 k=(4-2(-1/6)^3)/(-1/6)=-433/18

OpenStudy (callisto):

welcome!!~